# MyShadeQuoteStyleleftmargin=0pt, rightmargin=0pt, backgroundcolor=gray!25, linewidth=0pt, skipbelow=skipabove=

### v2

@@ Abstract.
Detecting vulnerabilities is vital for software security, yet deep learning-based vulnerability detectors (DLVD) face a data shortage, which limits their effectiveness. Data augmentation can potentially alleviate the data shortage, but augmenting vulnerable code is challenging and requires a generative solution that maintains vulnerability. Previous works have only focused on generating samples that contain single statements or specific types of vulnerabilities. Recently, large language models (LLMs) have been used to solve various code generation and comprehension tasks with inspiring results, especially when fused with retrieval augmented generation (RAG). Therefore, we propose VulScribeR, a novel LLM-based solution that leverages carefully curated prompt templates to augment vulnerable datasets. More specifically, we explore three strategies to augment both single and multi-statement vulnerabilities, with LLMs, namely Mutation, Injection, and Extension. Our extensive evaluation across three vulnerability datasets and DLVD models, using two LLMs, show that our approach beats two SOTA methods Vulgen and VGX, and Random Oversampling (ROS) by 27.48%, 27.93%, and 15.41% in f1-score with 5K generated vulnerable samples on average, and 53.84%, 54.10%, 69.90%, and 40.93% with 15K generated vulnerable samples. Our approach demonstrates its feasibility for large-scale data augmentation by generating 1K samples at as cheap as US\$ 1.88.
@@ 3. Methodology
  • Verifier The point of the Verifier is to have control over the quality of the generated code as there’s no guarantee that the LLM produces acceptable code. We use Joern’s (QwietAI, [n. d.]), a fuzzy C language parser, which is based on Antlr’s (Parr and Quong, 1995) C Parser, to filter out generated codes that contain severe syntax errors. The rationale behind using a fuzzy parser instead of a strict parser is that in data augmentation for various machine learning tasks, the generated data does not need to conform to the original data’s strict standards (DeVries and Taylor, 2017; Zhang et al., 2018; Guo et al., 2019; Li et al., 2022; Dong et al., 2023). Note that our goal in this study is to augment existing vulnerable datasets to help models capture the vulnerability patterns and generalize better, rather than generating high-quality vulnerable samples. Completely correct code is not necessary for training models (Wang et al., 2024) and data with subtle noise helps with the generalization of models (Bishop, 1995; Frenay and Verleysen, 2014; Hickey, 1996). Nevertheless, we acknowledge that more sophisticated methods can be used for the verification phase, but starting from a simple module (i.e. a parser) was necessary for evaluating whether using LLMs for vulnerability detection is even feasible. As a results, the fuzzy C Parser filters out 2%-13% of the generated data.
@@ 3.1.1. Mutation
The majority of previous works on program augmentation (Yu et al., 2022; Orvalho et al., 2022; Mao et al., 2024; Bui et al., 2021; Dong et al., 2023; Chakraborty et al., 2022), rely on code transforms that do not change the flow, semantics, and syntactical correctness of the program using program analysis. Variable name changing, replacing for with while loops and vice versa, and adding dead code are examples of this. However, program analysis is very time-consuming, and selecting the locations for transformation is challenging. Previous studies usually select the types and locations of transformation randomly without any program comprehension (Yu et al., 2022; Orvalho et al., 2022; Mao et al., 2024; Bui et al., 2021; Dong et al., 2023; Chakraborty et al., 2022). In Mutation strategy, we aim to augment vulnerable code samples by utilizing the program comprehension capability of LLM to mutate existing vulnerable code samples and let the LLM choose both the type of transformation and the potential statements to transform. In this way, we get around the program analysis and rely on LLM’s creativity to generate more suitable and diverse items. We design the prompt template as shown in Figure 2.
@@ 4.1. Research Questions
In RQ1, we evaluate the effectiveness of VulScribeR for improving the performance of DLVD models by augmenting vulnerable data, by comparing it to current SOTA approaches. In RQ2, we aim to investigate the effectiveness of our design for the Retriever component. In RQ3, we explore how the number of generated samples by our approach and SOTA methods affect the performance of DLVD models.
@@ 4.2. Datasets
In this study, we evaluate our approach on three different widely used vulnerability detection datasets (Nong et al., 2023, 2024; Yang et al., 2023; Fu and Tantithamthavorn, 2022; Zhou et al., 2019; Chakraborty et al., 2020), namely Devign (Zhou et al., 2019), Reveal (Chakraborty et al., 2020) and BigVul (Fan et al., 2020). Table 1 shows the details of these datasets. All of these datasets include C/C++ functions gathered from real-world projects.
@@ 4.2.1. Devign
Following previous studies (Nong et al., 2024, 2023) we use Devign as our primary training dataset as it has a better balance between the two categories, namely the vulnerable and clean samples (i.e. non-vulnerable samples) by following a previous study, VulGen (Nong et al., 2023). We also use the clean samples of this dataset as the clean inputs for the Retriever component.
@@ 4.7.2. Approach of RQ2
In this RQ, we investigate the usefulness of our proposed RAG for the Injection and Extension strategies. Specifically, we eliminate the Retriever component, and instead, we match the clean and vulnerable samples randomly to examine the effect of RAG (i.e. Injection w/o Retriever and Extension w/o Retriever).
Moreover, we also conduct an ablation study on the clustering phase of the Retriever component to examine the effect of diversity on the quality of the generated samples. To do so, we remove the clustering phase from the Retriever component and search for similar samples for each clean input item within all the vulnerable items instead of within each cluster. Then we take the top 5 retrieved items to end up with a similar number of $clean-vul$ pairs. In the sampling phase, we sort all the pairs based on their relevance score and start from the top. We denote the Injection strategy without clustering phase as Injection w/o Clustering and the Extension strategy without cluster as Extension w/o Clustering.
@@ 4.7.3. Approach of RQ3
Lastly, we investigate how the number of augmented samples impacts the performance of DLVD models, and how much performance can be gained by providing more data.
@@ 4.8. Implementation details
We use four 24GB Nvidia RTX 3090s for loading and prompting CodeQwen1.5, training, and testing LineVul, Devign, and Reveal models. We conducted a total of 616 experiments for all our RQs which cost more than 1500 GPU hours solely for training and testing DLVD models (excluding the time spent generating the samples).
@@ 5.1. RQ1: How effective is VulScribeR compared to SOTA approaches?
Table 2. Comparison of VulScribeR’s strategies with baselines using ChatGPT and CodeQwen when augmenting 5K Samples. The cells with larger values (better performance) compared to NoAug are highlighted darker.
@@
As discussed in Section 2, SOTA approaches like VGX and VulGen only focus on single-statement that limits the diversity of generated vulnerabilities. However, our approaches do not have this limitation and are able to generate more diverse vulnerable samples. To examine this, we measure the diversity of the augmented vulnerable samples by our approaches and VGX and VulGen. By following previous studies, we first apply principle component analysis (PCA) on CodeBERT’s embeddings of the generated vulnerable samples and reduce the dimension to three. Next, we calculate the histogram using 10 bins, from which we calculate the entropy of the vectors. Higher entropy values indicate greater diversity among the vulnerable samples, while lower values suggest a more concentrated distribution with similar samples. As depicted in Figure 5 the entropy for Mutation, Injection, and Extension are 4.79, 4.5, and 4.62 for ChatGPT, and 4.66, 4.45, 4.42 for CodeQwen, while for Vulgen and VGX, it is 4.42 and 4.31 respectively. The results demonstrate that our approaches have a better potential for generating more diverse samples as we beat them while using only one of the datasets they used for vulnerability mining. VulScribeR produces more diverse vulnerable samples compared to SOTA approaches VXG and VulGen.
@@ 5.2. RQ2: How does RAG contribute to VulScribeR?
Table 3. Results of Ablation Studies on the Retriever and Clustering for Injection.
@@
Table 4. Results of Ablation Studies on the Retriever and Clustering for Extension.
@@
Table 5. Impact of Generated Samples on Improving DVLD models’ Performance at 5K, 10K, and 15K
@@
Figure 6. The impact of the number of augmented samples on the effectiveness of DLVD models across the studied datasets and LLMs.

### v3

@@ Abstract.
Detecting vulnerabilities is vital for software security, yet deep learning-based vulnerability detectors (DLVD) face a data shortage, which limits their effectiveness. Data augmentation can potentially alleviate the data shortage, but augmenting vulnerable code is challenging and requires a generative solution that maintains vulnerability. Previous works have only focused on generating samples that contain single statements or specific types of vulnerabilities. Recently, large language models (LLMs) have been used to solve various code generation and comprehension tasks with inspiring results, especially when fused with retrieval augmented generation (RAG). Therefore, we propose VulScribeR, a novel LLM-based solution that leverages carefully curated prompt templates to augment vulnerable datasets. More specifically, we explore three strategies to augment both single and multi-statement vulnerabilities, with LLMs, namely Mutation, Injection, and Extension. Our extensive evaluation across four vulnerability datasets and DLVD models, using three LLMs, show that our approach beats two SOTA methods Vulgen and VGX, and Random Oversampling (ROS) by 27.48%, 27.93%, and 15.41% in f1-score with 5K generated vulnerable samples on average, and 53.84%, 54.10%, 69.90%, and 40.93% with 15K generated vulnerable samples. Our approach demonstrates its feasibility for large-scale data augmentation by generating 1K samples at as cheap as US\$ 1.88.
@@ 1. Introduction
It is worth noting that recent works have also explored vulnerability detection using LLMs (Du et al., 2024; Lu et al., 2024; Ding et al., 2024; Lin and Mohaisen, 2025), which pose as competitors to DLVD models. However, adopting LLM-based vulnerability detection methods poses various challenges and limitations that make DLVD models still highly relevant: 1) Data Privacy Concerns: in the industry, there are privacy concerns regarding the use of LLMs for vulnerability detection and essentially exposing private commercial code to LLMs; 2) Limited Robustness to Domain-specifc Vulnerabilities: LLMs require robust datasets to be used alongside various techniques, thus they usually struggle with project and domain-specifc vulnerabilities not present in the training data; 3) High Computational Costs: LLMs are resource-intensive to run while VulScribeR uses LLMs to augment vulnerable samples once and the trained DLVD models can be deployed while requiring significantly less compute and memory, making them easier to be deployed on-premise or in a constrained environment. Consequently, it is important to augment vulnerable data to improve the performance of DLVD models.
@@
  • RQ4: Can VulScribeR be effective on large, heavily-imbalanced datasets with complex real-world vulnerabilities?
    Results: on PrimeVul, both Injection and Extension outperform the baselines and Mutation by a considerable margin, while Extension provides a slightly higher performance compared to Injection. For instance, Extension outperforms NoAug, Vulgen, VGX, and ROS by 22.80%, 30.83%, 25.24%, and 3.87% on average F1-score.
@@ 3. Methodology
  • Verifier The goal of the Verifier is to have control over the quality of the generated code, as there’s no guarantee that the LLM produces compilable code. We use Joern’s (QwietAI, [n. d.]), a fuzzy C language parser, which is based on Antlr’s (Parr and Quong, 1995) C Parser, to filter out generated codes that contain severe syntax errors. The rationale behind using a fuzzy parser instead of a strict parser is that in data augmentation for various machine learning tasks, the generated data does not need to conform to the original data’s strict standards (DeVries and Taylor, 2017; Zhang et al., 2018; Guo et al., 2019; Li et al., 2022; Dong et al., 2025; Olsson et al., 2021; Yun et al., 2019). Note that our goal in this study is to augment existing vulnerable datasets to help models capture the vulnerability patterns and generalize better, rather than generating high-quality vulnerable samples. As such, completely correct code is not necessary for training models (Wang et al., 2024), and augmented code can improve the performance of the models even when they slightly break the source code syntax (Li et al., 2022; Dong et al., 2025); in fact, data with subtle noise helps with the generalization of models (Bishop, 1995; Frenay and Verleysen, 2014; Hickey, 1996). We use a parser instead of a compiler as we wish to keep the post-processing lightweight, as compilation needs significant efforts, such as identifying the dependencies and resolving type errors, which are unavailable in the original datasets, as they only contain functions. Nevertheless, we acknowledge that more sophisticated methods can be used for the verification phase, but starting from a simple module (i.e, a parser) was necessary for evaluating whether using LLMs for vulnerability detection is even feasible. As a result, the fuzzy C Parser filters out 2%-13% of the generated data.
@@ 3.1.1. Mutation
The majority of previous works on program augmentation (Yu et al., 2022; Orvalho et al., 2022; Mao et al., 2024; Bui et al., 2021; Dong et al., 2025; Chakraborty et al., 2022), rely on code transforms that do not change the flow, semantics, and syntactical correctness of the program using program analysis. Variable name changing, replacing for with while loops and vice versa, and adding dead code are examples of this. However, program analysis is very time-consuming, and selecting the locations for transformation is challenging. Previous studies usually select the types and locations of transformation randomly without any program comprehension (Yu et al., 2022; Orvalho et al., 2022; Mao et al., 2024; Bui et al., 2021; Dong et al., 2025; Chakraborty et al., 2022). In Mutation strategy, we aim to augment vulnerable code samples by utilizing the program comprehension capability of LLM to mutate existing vulnerable code samples and let the LLM choose both the type of transformation and the potential statements to transform. In this way, we get around the program analysis and rely on LLM’s creativity to generate more suitable and diverse items. We design the prompt template as shown in Figure 2.
@@ 4.1. Research Questions
  • RQ4: Can VulScribeR be effective on large, heavily-imbalanced datasets with complex real-world vulnerabilities?
In RQ1, we evaluate the effectiveness of VulScribeR for improving the performance of DLVD models by augmenting vulnerable data, by comparing it to current SOTA approaches. In RQ2, we aim to investigate the effectiveness of our design for the Retriever component. In RQ3, we explore how the number of generated samples by our approach and SOTA methods affects the performance of DLVD models. In RQ4, we investigate the usefulness of VulScribeR on a new large, high-quality, heavily-imbalanced dataset with real-world vulnerabilities.
@@
$\textbf{PrimeVul}_{\textit{Train}}$ 4802 169345 174147 1:35.3
$\textbf{PrimeVul}_{\textit{Validation}}$ 593 23347 23940 1:39.4
$\textbf{PrimeVul}_{\textit{Test}}$ 549 24232 18864 1:44.1
@@ 4.2. Datasets
In this study, we use three different widely used vulnerability detection datasets (Nong et al., 2023, 2024; Yang et al., 2023; Fu and Tantithamthavorn, 2022; Zhou et al., 2019; Chakraborty et al., 2020), namely Devign (Zhou et al., 2019), Reveal (Chakraborty et al., 2020), and BigVul (Fan et al., 2020), to evaluate our approach. Additionally, we evaluate our method with PrimeVul (Ding et al., 2024), which is a new large dataset with less label noise, in RQ4. Table 1 shows the details of these datasets. All of these datasets include C/C++ functions gathered from real-world projects. For Devign, Reveal, BigVul, and the extra clean data (Lin et al., 2019) used for retaining the ratio, we used the cleaned version from VGX (Nong et al., 2024), which had duplicates removed between the datasets. For PrimeVul, we used the SHA2 hash to remove samples available in Devign and the extra cleans (Lin et al., 2019) from all three sets of the PrimeVul dataset.
@@ 4.2.1. Devign
Following previous studies (Nong et al., 2024, 2023), we use Devign as our primary training dataset as it offers a better balance between the two categories, namely the vulnerable and clean samples (i.e., non-vulnerable samples) by following a previous study, VulGen (Nong et al., 2023). We also use the clean samples of this dataset as the clean inputs for the Retriever component.
@@ 4.2.3. Reveal
4.2.4. PrimeVul
Unlike VGX (Nong et al., 2024) and VulGen (Nong et al., 2023), which use Devign and BigVul datasets, which are considerably noisy (Croft et al., 2023; Risse and Böhme, 2024; Ding et al., 2024) (i.e., some vulnerable samples in these datasets are clean), we also use PrimeVul (Ding et al., 2024) in RQ4 with the settings of RQ1 with some changes (see more details in Section 5.4). PrimeVul is the largest dataset used in our study, and one of the most imbalanced datasets available. We use all three sets of Primevul in RQ4, with the train set replacing Devign as the main training set as well as BigVul’s train set for the retrieved vulnerable samples, and the validation and testing sets replacing those of BigVul.
@@ 4.5. Base LLMs
Additionally, we use GPT4o-mini (OpenAI, 2024) for our RQ4, which is as capable as ChatGPT3.5, but more affordable. Similar to ChatGPT3.5, we set the temperature parameter to 0.5, but we did not change any other hyperparameters.
@@ 4.7.2. Approach of RQ2
In this RQ, we investigate the usefulness of our proposed RAG for the Injection and Extension strategies. Specifically, we eliminate the Retriever component, and instead, we match the clean and vulnerable samples randomly to examine the effect of RAG (i.e. Injection w/o Retriever and Extension w/o Retriever). Moreover, we also conduct an ablation study on the clustering phase of the Retriever component to examine the effect of diversity on the quality of the generated samples. To do so, we remove the clustering phase from the Retriever component and search for similar samples for each clean input item within all the vulnerable items instead of within each cluster. Then we take the top 5 retrieved items to end up with a similar number of $clean-vul$ pairs. In the sampling phase, we sort all the pairs based on their relevance score and start from the top. We denote the Injection strategy without clustering phase as Injection w/o Clustering and the Extension strategy without clustering as Extension w/o Clustering.
@@ 4.7.3. Approach of RQ3
In this RQ, we investigate how the number of augmented samples impacts the performance of DLVD models, and how much performance can be gained by providing more data.
@@
4.7.4. Approach of RQ4
Lastly, we propose a slightly different setting than RQ1 for our approach to work with heavily imbalanced large datasets, and evaluate the effectiveness of our strategies.
For this RQ, we utilized GPT4-omini and used PrimeVul’s trainset as the primary training set as well as the source for collecting vulnerable samples that are fed to the Retriever, and used Devign’s cleans as the source of cleans. Since the PrimeVul dataset does not contain line-level information of vulnerabilities, we collected the vulnerable lines by computing the diff between paired clean and vulnerable samples and extracting the vulnerable lines.
It is worth mentioning that PrimeVul is a challenging dataset with a highly imbalanced ratio between vulnerable and clean samples (1:35) for training DLVD models (Ding et al., 2024). To make training possible, we balanced the dataset by randomly removing clean samples to achieve a ratio of 1:15 (larger ratios made training less effective on some models). Then we followed the settings of RQ1 for each strategy, and collected 5K new samples. To conduct a fair comparison between the methods, we used ROS to randomly oversample the original vulnerable data to get to a balanced dataset (i.e., 1:1 ratio) for training DLVD. Note that we do not change the ratio of the validation and testing sets.
@@ 4.8. Implementation details
We used a machine with four 24GB Nvidia RTX 3090s for our experiments. We conducted a total of 781 experiments for all our RQs, which cost close to 2000 GPU hours solely for training and testing DLVD models (excluding the time spent generating the samples).
@@ 5.1. RQ1: How effective is VulScribeR compared to SOTA approaches?
Table 2. Comparison of VulScribeR’s strategies with baselines using ChatGPT and CodeQwen when augmenting 5K samples across the studied DLVD models (i.e., Devign, Reveal, and Linevul). The cells with larger values (better performance) compared to NoAug are highlighted darker.
@@
As discussed in Section 2, SOTA approaches like VGX and VulGen only focus on single-statement that limits the diversity of generated vulnerabilities. However, our approaches do not have this limitation and are able to generate more diverse vulnerable samples. To examine this, we measure the diversity of the augmented vulnerable samples by our approaches and VGX and VulGen. By following previous studies, we first apply principle component analysis (PCA) on CodeBERT’s embeddings of the generated vulnerable samples and reduce the dimension to three. Next, we calculate the histogram using 10 bins, from which we calculate the entropy of the vectors. Higher entropy values indicate greater diversity among the vulnerable samples, while lower values suggest a more concentrated distribution with similar samples. As depicted in Figure 5 the entropy for Mutation, Injection, and Extension are 4.79, 4.5, and 4.62 for ChatGPT, and 4.66, 4.45, 4.42 for CodeQwen, while for Vulgen and VGX, it is 4.42 and 4.31 respectively. The results demonstrate that our approaches have a better potential for generating more diverse samples as we beat them while using only one of the datasets they used for vulnerability mining. VulScribeR produces more diverse vulnerable samples compared to SOTA approaches VGX and VulGen.
@@ 5.2. RQ2: How does RAG contribute to VulScribeR?
Table 3. Results of Ablation Studies on the Retriever and Clustering for Injection across the studied DLVD models (i.e., Devign, Reveal, and Linevul).
@@
Table 4. Results of Ablation Studies on the Retriever and Clustering for Extension across the studied DLVD models (i.e., Devign, Reveal, and Linevul).
@@
Table 5. Impact of Generated Samples on Improving DLVD models’ Performance at 5K, 10K, and 15K
@@
Figure 6. The impact of the number of augmented samples on the effectiveness of DLVD models across the studied datasets and LLMs across the studied DLVD models (i.e., Devign, Reveal, and Linevul).
@@
5.4. RQ4: Can VulScribeR be effective on large, heavily-imbalanced datasets with complex real-world vulnerabilities?
Table 6. Comparison of VulScribeR’s strategies on the PrimeVul dataset with baselines using GPT4o-mini when augmenting 5K Samples across the studied DLVD models (i.e., Devign, Reveal, and Linevul). The cells with larger values (better performance) compared to NoAug are highlighted darker.
Strategy Devign Reveal Linevul
P R F1 P R F1 P R F1
NoAug 25.93 5.12 8.55 7.35 44.63 12.62 18.31 31.15 23.06
VulGen 4.94 59.76 9.11 5.29 59.27 9.72 19.39 28.23 22.99
VGX 5.61 53.66 10.15 5.44 50.73 9.82 19.77 27.69 23.06
ROS 7.45 52.44 13.04 7.58 43.66 13.07 18.48 30.05 22.89
Mutation 7.17 52.93 12.63 7.81 43.66 13.25 15.71 26.96 19.85
Injection 7.57 55.37 13.31 7.81 49.76 13.51 17.29 31.88 22.42
Extension 7.61 53.41 13.33 8.51 41.22 14.11 18.51 31.15 23.22
Extension and Injection outperform Mutation. Extension slightly outperforms Injection. Mutation continues to struggle against ROS. Table 6 presents the results of VulScribeR and baselines across different DLVDs. As observed, Extension and Injection beat Mutation in all cases. Extension and Injection perform close to each other, while Extension appears to be more effective. In terms of F1-score, Extension achieves an F1-score of 16.89%, outperforming Injection (F1-score 16.41%) and Mutation (F1-score 15.24%) by 2.88% and 10.78% on average, respectively. One possible reason is that the Extension strategy creates samples that are closer to the decision boundaries compared to Injection, with the reason lying in the source of the vulnerable samples and cleans. In RQ1 5.1, for the Injection, most of the context, being the clean sections, comes from the same dataset, while in this RQ, this is true for Extension.
Extension appears to be useful and always outperforms baselines in all instances, with Injection performing close to Extension and outperforming all baselines with two of the three DLVDs. Extension outperforms NoAug, Vulgen, VGX, and ROS by 14.54%, 21.14%, 17.73%, and 3.39% on average F1-score, respectively. As shown in Table 6, Extension strategy outperforms baselines (i.e., NoAug, VulGen, VGX, ROS) in all of the experimental instances in terms of F1-score. The Injection fails to increase the performance for LineVul compared to the baselines. More specifically, Extension beats the baselines: NoAug, Vulgen, VGX, and ROS by 14.54%, 21.14%, 17.73%, and 3.39% on average F1-score, while the Injection strategy beats the baselines: NoAug, Vulgen, VGX, and ROS by 11.33%, 17.74%, 14.43%, and 0.49% on average F1-score, respectively.
As observed, Mutation only beats all the baselines with Reveal. Similar to RQ1 5.1, Mutation does not have a dominating advantage over ROS, and ROS beats Mutation in the two other settings. ROS beats the baseline, Vulgen, VGX, and Mutation by 10.78%, 17.17%, 13.97%, and 7.15% in terms of F1-score on average.
Extension outperforms the baselines and the Mutation by a large margin, while Injection provides a slightly lower performance compared to the Extension, but beats all the baselines with two DLVDs. Extension outperforms NoAug, Vulgen, VGX, and ROS by 14.54%, 21.14%, 17.73%, and 3.39% on average F1-score.
@@ 6. Discussion
Table 7. Impact of mixing our strategies’ generated samples on Improving DLVD models’ performance with a total of 5K augmented samples. Each involved strategy provides an equal number of samples. The cells with larger values (better performance) compared to NoAug are highlighted darker.
ChatGPT 3.5 Turbo CodeQwen1.5-7B-Chat
Strategy Devign Reveal Linevul Devign Reveal Linevul
P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
Reveal Mutation 10.98 53.44 18.22 12.23 68.46 20.75 7.80 22.62 11.60 10.96 61.64 18.60 15.01 57.96 23.85 9.48 30.43 14.45
M + I 11.89 66.41 20.17 13.74 70.39 22.99 11.46 47.14 18.43 11.78 58.02 19.58 13.14 72.74 22.25 9.35 51.41 15.82
M + E 11.23 58.02 18.82 14.14 59.89 22.87 9.45 43.23 15.50 11.92 55.19 19.60 13.20 65.98 22.00 7.85 42.51 13.26
I + E 13.39 47.71 20.91 15.26 59.71 24.31 22.07 9.01 12.80 14.15 46.98 21.75 14.92 62.24 24.06 10.34 18.80 13.35
M + I + E 11.12 46.98 17.98 9.67 53.26 16.36 10.55 47.14 17.25 11.07 57.60 18.57 13.34 75.09 22.65 40.33 9.47 15.34
Bigvul Mutation 7.80 63.91 13.90 7.63 67.73 13.72 6.78 74.14 12.42 7.56 57.23 13.36 9.37 53.26 15.94 6.65 78.31 12.25
M + I 7.28 66.93 13.14 8.79 66.45 15.53 7.41 72.75 13.45 7.30 58.82 12.99 8.31 64.71 14.72 7.03 83.60 12.97
M + E 7.24 61.21 12.94 8.68 57.55 15.09 6.94 81.65 12.78 7.45 61.84 13.30 8.27 58.51 14.49 6.64 82.76 12.30
I + E 9.41 35.29 16.46 10.83 39.90 17.04 8.53 41.80 14.16 11.11 31.96 16.49 10.42 38.47 16.40 7.46 34.01 12.24
M + I + E 7.14 42.77 15.20 9.25 53.42 15.77 7.38 79.61 13.50 7.45 54.85 13.12 9.23 52.78 15.71 6.83 78.59 12.57
Table 8. Impact of using random oversampling for clean samples instead of adding additional clean samples on our strategies and the baselines. Starred rows indicate using oversampled cleans. Bold F1-scores show the best performance between the two settings for each augmentation strategy.
ChatGPT 3.5 Turbo CodeQwen1.5-7B-Chat
Strategy Devign Reveal Linevul Devign Reveal Linevul
P R F1 P R F1 P R F1 P R F1 P R F1 P R F1
Reveal ROS 13.28 39.93 19.93 14.35 54.52 22.72 9.88 22.71 13.77 13.28 39.93 19.93 14.35 54.52 22.72 9.88 22.71 13.77
VulGen* 8.56 64.29 15.10 5.64 78.38 10.52 5.37 100.00 10.20 8.56 64.29 15.10 5.64 78.38 10.52 5.37 100.00 10.20
VGX* 8.86 65.98 15.62 5.74 74.40 10.66 5.38 99.91 10.21 8.86 65.98 15.62 5.74 74.40 10.66 5.38 99.91 10.21
Mutation* 9.05 59.95 15.72 9.90 73.34 17.45 4.66 44.77 8.44 9.07 66.77 15.98 6.25 86.17 11.65 5.66 70.12 10.47
Injection* 11.16 63.21 18.97 13.07 69.24 21.99 11.69 27.25 16.36 10.33 48.01 17.00 7.90 65.66 14.10 11.35 22.34 15.05
Extension* 10.51 56.82 17.73 12.52 54.46 20.36 12.33 30.43 17.55 10.24 46.62 16.80 7.23 53.74 12.74 6.83 60.58 12.28
Bigvul ROS 7.54 25.60 11.65 7.96 33.39 12.86 10.91 13.16 11.93 7.54 25.60 11.65 7.96 33.39 12.86 10.91 13.16 11.93
VulGen* 5.72 72.02 10.60 5.77 85.21 10.81 6.76 50.51 11.92 5.72 72.02 10.60 5.77 85.21 10.81 6.76 50.51 11.92
VGX* 5.82 69.16 10.74 5.75 77.27 10.71 6.98 49.03 12.22 5.82 69.16 10.74 5.75 77.27 10.71 6.98 49.03 12.22
Mutation* 6.33 70.75 11.63 6.39 78.22 11.82 5.96 83.50 11.12 6.20 82.51 11.53 6.15 85.69 11.48 6.08 92.22 11.40
Injection* 7.24 54.21 12.77 8.86 59.46 15.42 12.82 23.45 16.57 7.46 40.38 12.59 7.90 71.54 14.23 8.61 24.37 12.72
Extension* 7.76 47.85 13.35 7.58 51.67 13.22 11.75 24.28 15.84 8.48 41.18 14.06 6.86 52.15 12.12 6.67 71.73 12.21
6.1. Ensemble of Strategies
Since our strategies appeared to be useful, we decided to investigate whether a mix of strategies would perform better than using a single strategy. For such a comparison, we used the results of RQ1, and added three settings for using two strategies, with each contributing 2.5K samples, amounting to the same 5K extra samples, and an all-strategy setting where one third of the samples came from each of the strategies. Table 7 shows the result of this experiment. As observed, the combination does not always outperform the single strategy. In half (6 out of 12) of the settings, a mixed strategy works worse than a single strategy. Specifically, a mix of all three (M+I+E) and a combination of Mutation and Extension always fail to beat all of the single strategies. The only promising combination is the mix of Injection and Extension (I+E), which achieves the best performance in 4 out of 12 settings.
The best ensemble setting appears to be a combination of Injection and Extension settings, yet this setting appears to be less useful than only using Injection. Yet, since Extension appears to beat Injection in some settings (e.g. RQ4), we recommend using Injection and Extension strategies and a mix of these to get to achieve a higher performance.
6.2. Impact Analysis of the Extra Clean Samples
Following VGX and Vulgen, a proportional number of extra clean samples are added to the dataset to keep the ratio of vulnerable and clean samples in RQs 1-3. To examine the effect of the extra clean samples, we did an additional study alongside the results of RQ1, where instead of adding extra clean samples, we randomly oversampled the original clean samples to keep the ratio. Aside from our strategies, we also included VGX and Vulgen to see if we get a different result. As observed in Table 8, VGX and Vulgen get better performance in 10 out of 12 settings, while this is in all cases of Mutation, and 9 out of 12 settings for Injection and Extension. Specifically, Vulgen, VGX, Mutation, Injection, and Extension gain 22.75%, 20.07%, 28.4%, 12.76%, and 18.48% on average F1-score, respectively. Hence, Mutation appears to be the most dependent on the addition of extra clean samples, and Injection appears to be the least dependent method.
Compared to ROS (of RQ1), where only extra clean samples are added and vulnerable samples are oversampled, Vulgen, VGX, and Mutation drop 21.46%, 20.42%, and 17.21%, on average F1-score, respectively. Injection and Extension gain an improvement of 5.12% and 0.27% in terms of the average F1-score. This shows that the effect of vulnerability augmentation is more important for achieving a higher performance than adding extra clean samples for Injection and Extension, while the extra clean samples play a bigger role for Vulgen, VGX, and Mutation strategies.